How much energy is absorbed by a 3g substance that has a specific heat capacity of 0.7 J/gC and undergoes a temperature increase of 2 degrees?
1
Expert's answer
2020-06-04T10:38:08-0400

Q = m*q*"\Delta"T = 3*0.7*2 = 4.2 J